OpenStudy (anonymous):

Complete the ordered pair (14, ) for the equation x - 5y = -21. (14, ) is a solution.

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

thank you

Directrix (directrix):

I think the above answer is incorrect and that the answer should be y = 7. x - 5y = -21 14 - 5y = -21 14 -14 -5y = -21 - 14 -5y = -35 y = 7

OpenStudy (anonymous):

-I meant 7 but i write - when i write

OpenStudy (anonymous):

Thank you to the both of you

OpenStudy (anonymous):

-welcome